# PCR-99: A Practical Method for Point Cloud Registration with 99% Outliers

### v2

@@ IV OVERVIEW OF PCR-99
1 $L\leftarrow\mathbf{0}_{n\times n};$ for every possible pair $(i,j)$ do
@@ 3 $S\leftarrow\mathbf{0}_{1\times n};$ for $i=1,2,3,\cdots,n$ do
5 By sorting $S$ in a descending order, obtain the ranking numbers $r_{i}$ for $i=1,2,\cdots,n$; $\mathcal{I}_{\mathrm{largest}}\leftarrow\{\};\ n_{\mathrm{hypothesis}}\leftarrow 0;$ for $r_{\mathrm{sum}}=6,7,\cdots,3n-3$  do
6       $\left(r_{i}\right)_{\mathrm{L}}\leftarrow\max(1,r_{\mathrm{sum}}-2n+1);$ $\left(r_{i}\right)_{\mathrm{U}}\leftarrow\mathrm{floor}\left((r_{\mathrm{sum}}-3)/3\right);$ for $r_{i}=\left(r_{i}\right)_{\mathrm{L}},\cdots,\left(r_{i}\right)_{\mathrm{U}}$ do
7             $\left(r_{j}\right)_{\mathrm{L}}=\max(r_{i}+1,r_{\mathrm{sum}}-r_{i}-n);$ $\left(r_{j}\right)_{\mathrm{U}}=\mathrm{floor}\left((r_{\mathrm{sum}}-r_{i}-1)/2\right);$ for $r_{j}=\left(r_{j}\right)_{\mathrm{L}},\cdots,\left(r_{j}\right)_{\mathrm{U}}$ do
8                   $r_{k}\leftarrow r_{\mathrm{sum}}-r_{i}-r_{j};$ Find $(i,j,k)$ that corresponds to $(r_{i},r_{j},r_{k})$; if $(i,j,k)$ does not pass all of (24)–(26)  then
@@ 9                         Continue;
10                  Compute $(s,\mathbf{R},\mathbf{t})$ from $(i,j,k)$ using [30]; $\mathcal{I}\leftarrow$ All points that fit well in $(s,\mathbf{R},\mathbf{t})$ under the given inlier threshold $\delta_{\mathrm{in}}$; if $n(\mathcal{I})>n(\mathcal{I}_{\mathrm{largest}})$ then
@@ 11                         $\mathcal{I}_{\mathrm{largest}}\leftarrow\mathcal{I};$
12                  $n_{\mathrm{hypothesis}}\leftarrow n_{\mathrm{hypothesis}}+1;$ if $n_{\mathrm{hypothesis}}$ is a multiple of 1000 then
@@
14                               Compute $(s,\mathbf{R},\mathbf{t})$ from $\mathcal{I}_{\mathrm{largest}}$ using [30]; return $(s,\mathbf{R},\mathbf{t})$
15
16
17
18
@@ V-C Prescreening Test (Step 3 in Section IV)
Unknown
scale
RANSIC [17] 80 out of 5500 80 out of 5500
ICOS [18] 11 out of 5500 9 out of 5500
@@
PCR-99 1 out of 5500 0 out of 5500
Known
scale
RANSIC [17] 103 out of 5500 95 out of 5500
ICOS [18] 17 out of 5500 16 out of 5500
DANIEL [19] 8 out of 5500 7 out of 5500
TriVoC [31] 4 out of 5500 1 out of 5500
VODRAC [20] 6 out of 5500 3 out of 5500
VOCRA [29] 26 out of 5500 23 out of 5500
PCR-99 (no S/O) 5 out of 5500 1 out of 5500
PCR-99 5 out of 5500 0 out of 5500
@@
Unknown
scale
RANSIC [17] 15
ICOS [18] 7.7
PCR-99 (no S/O) 1.5
Known
scale
RANSIC [17] 153
ICOS [18] 31
DANIEL [19] 7.3
TriVoC [31] 1.3
VODRAC [20] 2.7
VOCRA [29] 1.1
PCR-99 (no S/O) 8.5
@@ VII RESULTS
We compare our method, PCR-99, against six open-source state-of-the-art methods: RANSIC [17], ICOS [18], DANIEL [19], TriVoC [31], VODRAC [20] and VOCRA [29]. We chose these methods for benchmarking, as they are claimed to be fast and robust up to 99% outliers. We also compare our method with an alternative version that adopts random sampling (as in RANSAC [11]) instead of the method described in Section V-B. At 99% outlier ratio, RANSIC and ICOS sometimes take hours to find a good solution. To prevent excessive computation times, we limit the iterations to 100 seconds for all methods. For known-scale problems, we evaluate all of the aforementioned methods. For unknown-scale problems, we only evaluate PCR-99 (with or without sample ordering), RANSIC and ICOS, as the other four methods are not applicable. All experiments are conducted in MATLAB on a laptop with i7-7700HQ (2.8 GHz) and 16 GB RAM.
Figure 2: Evaluation results for unknown-scale problems: RANSIC and ICOS are not robust at 99% outlier ratio when the computation time is limited to 100 seconds. In contrast, PCR-99 (with or without sample ordering) is robust up to 99% outlier ratio. The sample ordering has a very small impact on the accuracy and robustness of PCR-99. Comparing the median times of the two versions of PCR-99, the random sampling approach is faster up to 98% outlier ratio, but slower at 99% outlier ratio.
@@ VII-B Known-Scale Registration
Figure 3: Evaluation results for known-scale problems: TriVoC, VODRAC, and PCR-99 (with or without sample ordering) are more robust than the other methods, producing fewer large rotation errors (e.g., larger than $10^{\circ}$). They exhibit similar levels of robustness across all outlier ratios. For PCR-99, the sample ordering boosts the speed significantly at high outlier ratios, without causing a noticeable change in the accuracy or robustness. Among all methods, PCR-99 is the fastest one at 99% outlier ratio.

### v3

@@ IV OVERVIEW OF PCR-99
1
[…]
3for every possible pair $(i,j)$ do
@@ 9       Obtain $\mathcal{A}$ using (8)–(10);
12
13By sorting $S$ in a descending order, obtain the ranking numbers $r_{i}$ for $i=1,2,\cdots,n$;
[…]
15for $r_{\mathrm{sum}}=6,7,\cdots,3n-3$  do
16       $\left(r_{i}\right)_{\mathrm{L}}\leftarrow\max(1,r_{\mathrm{sum}}-2n+1);$
17      $\left(r_{i}\right)_{\mathrm{U}}\leftarrow\mathrm{floor}\left((r_{\mathrm{sum}}-3)/3\right);$
18      for $r_{i}=\left(r_{i}\right)_{\mathrm{L}},\cdots,\left(r_{i}\right)_{\mathrm{U}}$ do
19
20            $\left(r_{j}\right)_{\mathrm{L}}=\max(r_{i}+1,r_{\mathrm{sum}}-r_{i}-n);$
21            $\left(r_{j}\right)_{\mathrm{U}}=\mathrm{floor}\left((r_{\mathrm{sum}}-r_{i}-1)/2\right);$
22            for $r_{j}=\left(r_{j}\right)_{\mathrm{L}},\cdots,\left(r_{j}\right)_{\mathrm{U}}$ do
23                   $r_{k}\leftarrow r_{\mathrm{sum}}-r_{i}-r_{j};$
24                  Find $(i,j,k)$ that corresponds to $(r_{i},r_{j},r_{k})$;
25                  if $(i,j,k)$ does not pass all of (24)–(26)  then
@@ 26                         Continue;
27
28                  Compute $(s,\mathbf{R},\mathbf{t})$ from $(i,j,k)$ using [30];
29                  $\mathcal{I}\leftarrow$ All points that fit well in $(s,\mathbf{R},\mathbf{t})$ under the given inlier threshold $\delta_{\mathrm{in}}$;
30                  if $n(\mathcal{I})>n(\mathcal{I}_{\mathrm{largest}})$ then
@@ 31                         $\mathcal{I}_{\mathrm{largest}}\leftarrow\mathcal{I};$
32
33                  $n_{\mathrm{hypothesis}}\leftarrow n_{\mathrm{hypothesis}}+1;$
34                  if $n_{\mathrm{hypothesis}}$ is a multiple of 1000 then
@@ 33                  $n_{\mathrm{hypothesis}}\leftarrow n_{\mathrm{hypothesis}}+1;$
36                               Compute $(s,\mathbf{R},\mathbf{t})$ from $\mathcal{I}_{\mathrm{largest}}$ using [30];
37                              return $(s,\mathbf{R},\mathbf{t})$
38
39
40
41
@@ V-C Prescreening Test (Step 3 in Section IV)
Unknown scale RANSIC [17] 79 out of 5500 79 out of 5500
ICOS [18] 12 out of 5500 11 out of 5500
@@
PCR-99 0 out of 5500 0 out of 5500
Known scale RANSIC [17] 81 out of 5500 78 out of 5500
ICOS [18] 24 out of 5500 22 out of 5500
DANIEL [19] 5 out of 5500 3 out of 5500
TriVoC [31] 2 out of 5500 0 out of 5500
VODRAC [20] 9 out of 5500 2 out of 5500
VOCRA [29] 16 out of 5500 16 out of 5500
PCR-99 (no S/O) 4 out of 5500 0 out of 5500
PCR-99 4 out of 5500 0 out of 5500
@@
Unknown scale RANSIC [17] 16
ICOS [18] 7.0
PCR-99 (no S/O) 1.6
Known scale RANSIC [17] 162
ICOS [18] 33
DANIEL [19] 8.0
TriVoC [31] 1.4
VODRAC [20] 3.0
VOCRA [29] 1.2
PCR-99 (no S/O) 9.3
@@ VII RESULTS
We compare our method, PCR-99, against six open-source state-of-the-art methods: RANSIC [17], ICOS [18], DANIEL [19], TriVoC [31], VODRAC [20] and VOCRA [29]. We chose these methods for benchmarking, as they are claimed to be fast and robust up to 99% outliers. RANSAC [11], FGR [25], GORE [10], GNC-TLS [24] and TEASER [16] are not included in the evaluation, as they are shown to be outperformed by the aforementioned methods (when parallelization is disabled for the fair runtime evaluation) [17, 18, 19, 31, 20, 29]. We also compare our method with an alternative version that adopts random sampling (as in RANSAC [11]) instead of the method described in Section V-B. At 99% outlier ratio, RANSIC and ICOS sometimes take hours to find a good solution. To prevent excessive computation times, we limit the iterations to 100 seconds for all methods. For known-scale problems, we evaluate all of the aforementioned methods. For unknown-scale problems, we only evaluate PCR-99 (with or without sample ordering), RANSIC and ICOS, as the other four methods are not applicable. All experiments are conducted in MATLAB on a laptop with i7-7700HQ (2.8 GHz) and 16 GB RAM.
Figure 2: Evaluation results for unknown-scale problems (500 Monte Carlo runs): RANSIC and ICOS are not robust at 99% outlier ratio when the computation time is limited to 100 seconds. In contrast, PCR-99 (with or without sample ordering) is robust up to 99% outlier ratio. The sample ordering has a very small impact on the accuracy and robustness of PCR-99. Comparing the median times of the two versions of PCR-99, the random sampling approach is faster up to 98% outlier ratio, but slower at 99% outlier ratio.
@@ VII-B Known-Scale Registration
Figure 3: Evaluation results for known-scale problems (500 Monte Carlo runs): TriVoC, VODRAC, and PCR-99 (with or without sample ordering) are more robust than the other methods, producing fewer large rotation errors (e.g., larger than $10^{\circ}$). They exhibit similar levels of robustness across all outlier ratios. For PCR-99, the sample ordering boosts the speed significantly at high outlier ratios, without causing a noticeable change in the accuracy or robustness. Among all methods, PCR-99 is the fastest one at 99% outlier ratio.